# Resona: Improving Context Copying in Linear Recurrence Models with Retrieval

### v1

@@ 1 Introduction
Corresponding author: Xinyu Wang (xinyu.wang5@mail.mcgill.ca), affiliated with McGill University, and Noah’s Ark Lab, Canada.
Improvements in building state-of-the-art LLMs (OpenAI, 2024; Grattafiori et al., 2024; Qwen, 2024; Gemma Team, 2024) through increased scale (Chung et al., 2024; Kaplan et al., 2020) and downstream tuning (Ouyang et al., 2022; Dubois et al., 2023) have enabled them to attain human-level performance on a number of complex tasks. One particular feature that has enabled such a feat is the use of in-context learning (Brown et al., 2020), where models can use user-provided content to provide a specific response tailed to that example. This relies on the Transformer (Vaswani et al., 2017) backbone that underlies many of these models, enabling for models to observe the complete past when generating content.
@@
However, a growing trend is the emergence of linear recurrent models (LRMs) (Gu et al., 2022; Peng et al., 2023; Orvieto et al., 2023; Gu and Dao, 2024), aimed to address a number of computational bottlenecks related to the use of attention (Bahdanau et al., 2015). Unlike Transformer-based models, LRMs do not operate over all previous parts of the input but instead compress it within a unified hidden representation/state much like original recurrent neural networks (RNNs) (Rumelhart, 1989; Hochreiter and Schmidhuber, 1997; Cho et al., 2014), enabling for more efficient inference. However, this unified representation has been shown to also act as an information bottleneck, where the number of potential elements in the vocabulary cannot be properly maintained within a state that does not grow with the possible combinations of tokens. This has raised questions regarding the ability of such models to adequately learn from input contexts (Jelassi et al., 2024; Park et al., 2024; Grazzi et al., 2024) and perform comparably with similar Transformer-based LLMs in such settings.
@@
Linear methods rely on fixed-size states to account limitations of traditional attention. A trade-off, however, is these fixed-size states cannot perfectly preserve all historical information, making exact retrieval challenging. This manifests itself in practice in various tasks such as language modeling; with the key-value associative memory system that underlies such methods, adding new key-value associations leads to accumulating retrieval errors that hinder performance. These errors can accumulate, whether it be data-independent (Gu et al., 2022; Orvieto et al., 2023) or dependent (Yang et al., 2024a; Gu and Dao, 2024; Peng et al., 2024; Zhang et al., 2024; Beck et al., 2017), and despite proposals to correct such errors (Yang et al., 2024b; 2025; Sun et al., 2024), these retain a fixed-size hidden state that cannot address the aforementioned limitation.
@@
In an effort to bridge this gap, we introduce Resona as a method to improve context-based learning in LRMs through retrieval, Resona (Figure 1). we introduce a retrieval mechanism that enables the information flow from the context to overcome bottlenecks in the hidden state, enabling the model to better use it for in-context learning. First, specific LRM layers within the backbone model are augmented with Resona, which consists of a mechanism that performs a contextual search on the input context, which we implement by chunking the context into passages and retrieving them based on the LRM state. This is followed by a knowledge-integration component that mixes the retrieved context back into the LRM output by directly modifying the output representation. This enables for information from the context to overcome the hidden state bottleneck, allowing for improved information flow from the context to the generated input. These processes (chunking, retrieval, and integration) are further parallelized with the main LRM, ensuring easy adaptation and improved in-context learning across a variety of models.
@@
Empirical results on a set of representative tasks, both synthetic and real-world, demonstrate that Resona significantly improves the abilities of LRMs to utilize context-specific information with minimal to no latency cost. We evaluate Resona on a diverse set of tasks, including synthetic retrieval and recall tasks, language modeling and question-answering tasks, across multiple scenarios such as pre-training and direct-finetuing. Our analysis demonstrates the effectiveness of using Resona for overall performance improvements as well as test-time model customizations, such as balancing the trade-off between efficiency and performance.
@@ 2 Related Works
Despite the vast improvements in constructing language models that solve real-world natural language tasks since the introduction of the Transformer (Vaswani et al., 2017), significant concerns remain regarding the efficiency and scalability of such models. While this has spurred interest in rendering them more efficient (Katharopoulos et al., 2020; Dao et al., 2022; Yang et al., 2024a), LRMs (Gu et al., 2022; Orvieto et al., 2023; Qin et al., 2023; Gu and Dao, 2024; Dao and Gu, 2024; Peng et al., 2023) have grown as a popular alternative due to highly efficient inference costs compared to attention-based alternatives while retaining the ability to be train on elements of a sequence in parallel, an issue with traditional recurrent models. Further attempts at leveraging advantages from both paradigms (Lieber et al., 2024; De et al., 2024; Dong et al., 2025) has also garnered interest. Nevertheless, these works fail to resolve exiting criticisms regarding the use of context and properly integrating such models with retrieval-based techniques.
Retrieval-Augmented Generation.
@@
RAG-based methods augment the input of a LM with passages retrieved from an outside source (Guu et al., 2020; Lewis et al., 2020). Such methods have significantly improved performance on knowledge-intensive tasks, where the amount of information necessary for strong performance is difficult to store explicitly within parametric knowledge (Roberts et al., 2020). Further improvements have come under the form of improved filtering of retrieved passages (Asai et al., 2024), robustness to irrelevant passages Yoran et al. (2024); Xu et al. (2024) or tuning of more components (Lin et al., 2024). However, RAG is not directly applicable for learning from contexts, as such methods do not search within the input-specific query but rather from an external database, nor do they account for the information bottleneck with LRMs. As such, we propose Resona as a method to overcome these limitations.
Do LRMs Use Context Effectively?
@@
Despite their practical benefits, questions have arose regarding the ability of LRMs to learn from input contexts (Akyürek et al., 2024). Jelassi et al. (2024) show that they struggle to directly copy information from contexts due to their fixed-sized latent state. Park et al. (2024) further show that they can struggle at retrieval-based (Arora et al., 2024) in-context learning, only solving such tasks through the addition of attention. Such observations have extended to real-world data, where LRMs have been shown to exhibit similar difficulties as Transformer LLMs (Huang, 2025; Liu et al., 2024) for long contexts (Ivgi et al., 2023; Hsieh et al., 2024). Accordingly, we introduce Resona as a potential solution by as a way of providing additional information flow paths from the context to the generated input, enabling better utilization of the context for problem-solving.
Memory-enhanced Transformer.
@@
Due to the quadratic complexity of self-attention, transformer models face significant computational challenges when processing long inputs. Numerous approaches have been proposed to enhance transformers for long sequential data, often incorporating memory mechanisms or retrieval modules.  Dai et al. (2019); Munkhdalai et al. (2024) segment long inputs into shorter sequences and process them recurrently.  Mohtashami and Jaggi (2023) appends landmark tokens to represent each block of input and uses group attention to select relevant information.  Borgeaud et al. (2022) enhances transformer performance with external data by leveraging a seperated retriever module.  Nunez et al. (2024) develop a Span-Expanded Attention for the hybridized attention model to retrieve the most relevant block and integrates it with the recent context for attention computation. However, it remains unclear whether and how retrieval-based modules can enhance the performance of generalized linear recurrent models such as GLA Yang et al. (2024a), Mamba Gu and Dao (2024); Dao and Gu (2024), RWKV Peng et al. (2023), and DeltaNet Yang et al. (2024b).
@@ 3 Resona
We introduce Resona (Figure 2 and Algorithm 1) as a framework to enhance the context-copying abilities of LRMs through retrieval, without sacrificing original performance and versatility. Our end-to-end training enables models to utilize the context as a retrieval base from which information can be extracted, enabling the model to overcome the fixed-size latent space bottleneck and integrate information from the context directly into the hidden state. This is in contrast to traditional LRMs, where the information from the context must flow through the hidden state.
@@
1:Model $\mathcal{M}$
2:Input: Input ${\bm{s}}\in\mathbb{R}^{T}$ and LRM hidden state ${\bm{h}}\in\mathbb{R}^{L\times T\times H}$, Output: ${\bm{y}}\in\mathbb{R}^{T\times D}$
3:Embed sequence ${\bm{s}}$ into ${\bm{x}}\in\mathbb{R}^{T\times D}$.
4:for $l\leftarrow 1$ to $L$ do
5:     if Layer $l$ is an Resona layer then
6:         ${\bm{h}}_{l},{\bm{y}}^{m}=\mathcal{M}_{l}^{\text{LRM}}\left({\bm{x}},{\bm{h}}_{l}\right)$
7:         ${\bm{M}}=\mathcal{M}_{l}^{\text{C-and-S}}({\bm{x}},{\bm{h}}_{l})$
8:         ${\bm{y}}^{r}=\mathcal{M}_{l}^{\text{KI}}\left({\bm{x}},{\bm{h}}_{l},{\bm{M}}\right)$
9:         $\bm{\alpha}=f({\bm{x}})$
10:         ${\bm{y}}=\bm{\alpha}\cdot{\bm{y}}^{m}+(1-\bm{\alpha}){\bm{y}}^{r}$
11:     else
12:         ${\bm{h}}_{l},{\bm{y}}=\mathcal{M}_{l}\left({\bm{x}},{\bm{h}}_{l}\right)$
13:     ${\bm{x}}={\bm{y}}$
14:return ${\bm{y}}$
@@ 3.1 Problem Formulation and Overview
Let $\mathbb{V}$ be a vocabulary of size $V$. A model $\mathcal{M}$ applies a function $f:\mathbb{V}^{*}\to\mathbb{V}^{*}$, which maps an input sequence of tokens to an output sequence ${\bm{y}}$. We also refer to the input ${\bm{x}}=[x_{1},\dots,x_{T}$] as the “prompt” to the model $\mathcal{M}$ and of the output sequence ${\bm{y}}=f({\bm{x}})$ as the generated “answer”. Furthermore, a sequence-to-token mapping is a function $g:\mathbb{V}^{*}\to\mathbb{V}$ that defines $f$ by auto-regressive inference. Namely, for every input sequence ${\bm{x}}\in\mathbb{V}^{*}$, $x_{i+j}$ is defined recursively as $x_{i+j}=g(x_{1},\dots,x_{i+j-1})$ and $f({\bm{x}})=(x_{i+1},x_{i+2},\dots)$.
A linear recurrent model (LRM) is defined by a state update rule $u:\mathcal{S}\times\mathbb{V}\to\mathcal{S}$ and an output function $r:\mathcal{S}\to\mathbb{V}$, where $\mathcal{S}$ is a finite set of states. Let $s_{0}\in\mathcal{S}$ be some initial state. Given some sequence ${\bm{x}}$ of length $L$, the state of the model at iteration $i$ is denoted by $S_{i}(x_{1},\dots,x_{i})$ and the output token is denoted by $R_{i}(x_{1},\dots,x_{i})$. These are defined recursively as:
@@
    ${\bm{y}}_{i}=R_{i}(x_{1},\dots,x_{i})=r\left(S_{i}(x_{1},\dots,x_{i})\right)$.
Information from ${\bm{x}}$ reaches ${\bm{y}}$ through a state ${\bm{h}}\in\mathbb{R}^{H}$ where $H$ is fixed and finite. Thus for increasing sequence length or information dense settings, such LRMs can struggle from the limited of ${\bm{h}}$. This can be formalized more explicitly as
Lemma 3.1.
If $\left|\mathcal{S}\right|>2^{H\times\mathrm{p}}$, with $\mathrm{p}$ the precision of ${\bm{h}}$, then ${\bm{h}}$ serves as an information bottleneck.
Proof.
Suppose $r$ applies a different function on every state $s\in\mathcal{S}$. Then ${\bm{h}}$ needs to have the capacity to model every possible state $s$. If $\left|\mathcal{S}\right|>2^{H\times\mathrm{p}}$, then by consequence there must be a subset $S^{\prime}\subseteq\mathcal{S}$ such that ${\bm{h}}$ represents at least 2 states $s$ and $s^{\prime}$ with the same encoding. Suppose that no information is lost. Then this means that for any update rule $u$ or $r$, applying these on $s$ and $s^{\prime}$ has the same result. However, this means that $s$ and $s^{\prime}$ are in fact the same state, hence there is no information loss by contradiction. ∎
Thus we observe that the LRM is directly limited by the size of its hidden state, which can be insufficient for modeling more complex problems unless it grows with the size of the possible set of states. Resona introduces two flexible components to overcome this constraint without sacrificing the primary benefits of LRMs (namely parallel training and inference time efficiency): 1) a contextual search operation that operates on the input to retrieve context-specific information and 2) a knowledge integration component that mixes the retrieved information with the LRM output.
@@
2:Input: Input ${\bm{x}}\in\mathbb{R}^{T\times D}$ and LRM hidden state ${\bm{h}}\in\mathbb{R}^{T\times H}$, Output: Attention mask ${\bm{M}}\in\mathbb{R}^{T\times T}$
3:Chunk ${\bm{x}}$ into ${\bm{x}}^{\prime}\in\mathbb{R}^{N\times U\times D}$
4:Use $\mathcal{C}$ to encode each context chunk ${\bm{x}}^{\prime}$ into context embeddings $\bar{\bm{C}}\in\mathbb{R}^{N\times E}$
5:Use $\mathcal{Q}$ to encode ${\bm{h}}$ into query embeddings $\bar{\bm{Q}}\in\mathbb{R}^{T\times E}$.
6:Compute chunk index sets for each $\bar{\bm{Q}}_{j}$: $\{I_{j}\}_{j=1}^{T}=\{\text{Top-}k(\bar{\bm{Q}}_{j},\bar{\bm{C}})\}_{j=1}^{T}$
7:With $\{I_{j}\}_{j=1}^{T}$, compute a mask ${\bm{M}}\in\mathbb{R}^{T\times T}$ such that ${\bm{M}}_{ji}=1\iff\left(i\in\{I_{j}\}\right)$.
@@
In order to retrieve relevant context, Resona implements contextual search as a chunk-and-search mechanism (Algorithm 2). The initial input ${\bm{x}}$ is first split into $N$ chunks, each consisting of $U$ tokens, to create ${\bm{x}}^{\prime}\in\mathbb{R}^{N\times U\times D}$. First, each of these chunks $C_{i}$ is encoded using a context encoder $\mathcal{C}$ into a context embedding $\bar{C}_{i}$. Simultaneously, the hidden state of the adjacent linear-recurrent layer ${\bm{h}}\in\mathbb{R}^{T\times H}$, is used to encode a number of queries into query embeddings $\bar{Q}_{1:T}$ using a query encoder $\mathcal{Q}$. For each query, we search for the top-$k$ similar contexts using a cosine-similarity search, which produces chunk indices that we can then use to retrieve the relevant input token positions. These are used to create a mask, which is passed to the Knowledge-Integration module.
@@
$\alpha$ can be computed on an input-dependent basis for each element of ${\bm{y}}$ or can alternatively be set as fixed hyper-parameter value. Because only the chunk-and-search design, each query attends to at most $kU$ elements from the initial input in $k$ contiguous blocks, enabling us to compute ${\bm{y}}^{r}$ efficiently using existing sparse attention mechanisms.
@@
2:Input: Input ${\bm{x}}\in\mathbb{R}^{T\times D}$ and LRM hidden state ${\bm{h}}\in\mathbb{R}^{T\times H}$, mask ${\bm{M}}\in\mathbb{R}^{T\times T}$, Output: ${\bm{y}}^{r}\in\mathbb{R}^{T\times D}$
3:From ${\bm{h}}$, compute queries ${\bm{Q}}\in\mathbb{R}^{T\times d_{k}}$ using $W_{Q}$. In parallel, compute ${\bm{K}},{\bm{V}}\in\mathbb{R}^{T\times d_{k}}$ from ${\bm{x}}$ with $W_{K}$, $W_{V}$.
4:Compute multi-head attention output ${\bm{O}}=\mathrm{CrossAttn}\left({\bm{Q}},{\bm{K}},{\bm{V}},{\bm{M}}\right)$, where ${\bm{O}}\in\mathbb{R}^{T\times d_{k}}$.
5:Project ${\bm{O}}$ using $W_{\text{out}}$ into ${\bm{y}}^{r}\in\mathbb{R}^{T\times D}$.

### v3

# Xinyu Wang   Linrui Ma11footnotemark: 1   Jerry Huang♡♢11footnotemark: 1   Peng Lu Prasanna Parthasarathi   Xiao-Wen Chang   Boxing Chen   Yufei Cui

@@ 1 Introduction
Improvements in building state-of-the-art large language models (LLMs) (OpenAI, 2024; Grattafiori et al., 2024; Qwen, 2024; Gemma Team, 2024) through increased scale (Chung et al., 2024; Kaplan et al., 2020) and downstream tuning (Ouyang et al., 2022; Dubois et al., 2023) have enabled them to attain human-level performance on a number of complex tasks. One feature that has enabled this is in-context learning (Brown et al., 2020), where models can use user-provided content to provide a specific response tailed to that example. This relies on the Transformer (Vaswani et al., 2017) backbone that underlies many of these models, enabling for models to observe the complete past when generating content.
@@
Recently, linear recurrent models (LRMs) (Gu et al., 2022; Peng et al., 2023; Orvieto et al., 2023; Gu and Dao, 2024; Yang et al., 2024b) have emerged as an alternative, aiming to address computational bottlenecks associated with attention mechanisms (Bahdanau et al., 2015). Unlike Transformers, LRMs do not operate over all previous parts of the input. Instead they compress prior context into a unified hidden representation/state with a recurrent structure, similar to original recurrent neural networks (RNNs) (Rumelhart, 1989; Hochreiter and Schmidhuber, 1997; Cho et al., 2014; Balduzzi and Ghifary, 2016; Lu et al., 2019; Martin and Cundy, 2018), enabling more efficient inference. However, this unified representation introduces an information bottleneck, as it limits the capacity to represent the full range of vocabulary elements within a fixed-size state that does not scale with the combinatorial complexity of token sequences. This has raised questions about the ability of LRMs to effectively learn from input contexts (Jelassi et al., 2024; Park et al., 2024; Grazzi et al., 2024; Lu et al., 2023) and perform comparably to Transformer-based LLMs in such settings.
@@
Linear methods rely on fixed-size states to address limitations of traditional attention. A trade-off, however, is that these fixed-size states cannot perfectly preserve all historical information, making exact retrieval challenging. This manifests itself in practice in various tasks such as language modeling; with the key-value associative memory system that underlies such methods, adding new key-value associations leads to accumulating retrieval errors that hinder performance. These errors may build up in a variety of data-independent (Gu et al., 2022; Orvieto et al., 2023) or dependent (Yang et al., 2024a; Gu and Dao, 2024; Peng et al., 2024; Zhang et al., 2024; Beck et al., 2017) manners. While recent works have proposed strategies to mitigate such errors (Yang et al., 2024b; 2025; Sun et al., 2024), they retain a fixed-size hidden state that remains a fundamental constraint.
@@
In an effort to bridge this gap, we propose Resona, a retrieval-based method designed to improve context-based learning in LRMs (Figure 1). By introducing a retrieval mechanism that facilitates information flow from the context, Resonamitigates the hidden state bottleneck and enables more effective in-context learning. Specifically, we augment LRM layers within the backbone model with a contextual search component. The input context is first chunked into passages, which are then retrieved based on the current LRM state. A knowledge integration module subsequently incorporates the retrieved passages into the model’s output by directly modifying its representation. This architecture allows previous context to bypass the fixed-size hidden state constraint, improving information flow from context to generation.These processes (chunking, retrieval, and integration) are parallelized with the main LRM, ensuring easy adaptation and improved in-context learning across a variety of models.
@@
Empirical results on a range of representative tasks, spanning both synthetic and real-world data, demonstrate that Resona significantly improves the ability of LRMs to utilize context-specific information with minimal or no latency overhead. We evaluate Resona on a diverse set of tasks, including synthetic retrieval and recall tasks, language modeling and question-answering tasks, across multiple scenarios such as pre-training and direct fine-tuning. Our analysis demonstrates the effectiveness of using Resona for overall performance improvements as well as test-time model customizations, such as balancing the trade-off between efficiency and performance.
@@ 2 Related Works
Despite vast improvements in building language models that solve real-world natural language tasks since the introduction of the Transformer (Vaswani et al., 2017), significant concerns remain about their efficiency and scalability. While this has spurred interest in rendering them more efficient (Katharopoulos et al., 2020; Dao et al., 2022; Yang et al., 2024a), LRMs (Gu et al., 2022; Orvieto et al., 2023; Qin et al., 2023; Gu and Dao, 2024; Dao and Gu, 2024; Peng et al., 2023; Sun et al., 2023; Lu et al., 2025; Yang et al., 2024b) have grown as a popular alternative due to highly efficient inference costs compared to attention-based alternatives while retaining the ability to be train on elements of a sequence in parallel, an issue with traditional recurrent models. Further attempts at leveraging advantages from both paradigms (Lieber et al., 2024; De et al., 2024; Dong et al., 2025) have also garnered interest, leading to further exploration of similar models.
Retrieval-Augmented Generation (RAG).
@@
RAG-based methods augment the input of an LM with passages retrieved from an outside source (Guu et al., 2020; Lewis et al., 2020). Such methods have significantly improved performance on knowledge-intensive tasks, where it is difficult to store the information required for strong performance explicitly within the model’s parametric knowledge (Roberts et al., 2020). Further improvements have come under the form of improved filtering of retrieved passages (Asai et al., 2024), robustness to irrelevant passages (Yoran et al., 2024; Xu et al., 2024) or tuning of more components (Lin et al., 2024). However, RAG is not directly applicable to learning from contexts, as such methods do not search within the input-specific query but rather from an external database, leading to a critical failure point of such methods.
Linear Recurrent Models and Context Usage.
@@
Despite their practical benefits, questions have arose regarding the ability of LRMs to learn from input contexts (Akyürek et al., 2024). Jelassi et al. (2024) show that they struggle to directly copy information from contexts due to their fixed-sized latent state. Park et al. (2024) further show that they can struggle at retrieval-based (Arora et al., 2024) in-context learning, only solving such tasks through the addition of attention. Such observations have extended to real-world data, where LRMs have been shown to exhibit similar difficulties as Transformer LLMs (Wang et al., 2024; Huang, 2025; Liu et al., 2024; Huang et al., 2025) for long contexts (Ivgi et al., 2023; Hsieh et al., 2024). Accordingly, we introduce Resona as a potential solution that provides additional information flow paths from the context to the generated input, enabling better utilization of the context for problem-solving.
Memory-enhanced Transformers.
@@
Due to the quadratic complexity of self-attention with respect to the length of a sequence, Transformer models face significant computational challenges when processing long inputs. Numerous approaches have been proposed to enhance Transformers for long sequential data, both to reduce the time/space complexity of the models as well as to improve performance. Dai et al. (2019); Munkhdalai et al. (2024) segment long inputs into shorter sequences and process them recurrently. Mohtashami and Jaggi (2023) append landmark tokens to represent each block of input and uses group attention to select relevant information. Borgeaud et al. (2022) enhance Transformer performance with external data by leveraging a separated retriever module. Nunez et al. (2024) develop a Span-Expanded Attention for the hybridized attention model to retrieve the most relevant block and integrate it with the recent context for attention computation. However, it remains unclear whether and how retrieval-based modules can enhance the performance of generalized linear recurrent models such as GLA (Mao, 2022; Yang et al., 2024a; Lu et al., 2025), Mamba (Gu and Dao, 2024; Dao and Gu, 2024), RWKV (Peng et al., 2023), and DeltaNet (Schlag et al., 2021; Yang et al., 2024b).
@@ 3 Resona
We introduce Resona (Figure 2 and Algorithm 1) as a framework to enhance the context-copying ability of LRMs through retrieval, without sacrificing original performance and versatility. Our end-to-end training enables models to utilize the context as a retrieval base from which information can be extracted. This helps the model to first overcome the fixed-size latent space bottleneck by integrating information from the context directly into the hidden state. This is in contrast to traditional LRMs, where the information from the context must flow through the hidden state.
@@
1:Model $M$
[…]
3:Output: ${\bm{Y}}\in\mathbb{R}^{T\times D}$
4:Embed sequence ${\bm{s}}$ into ${\bm{X}}\in\mathbb{R}^{T\times D}$.
5:for $l\leftarrow 1$ to $L$ do
[…]
7:     if Layer $l$ is an Resona layer then
8:         $({\bm{H}},{\bm{Y}}^{m})=M_{l}^{\text{LRM}}\left({\bm{X}},{\bm{H}}\right)$
9:         $\triangleright$ LRM Output
10:         ${\bm{M}}=M_{l}^{\text{C-and-S}}({\bm{X}},{\bm{H}})$
11:         $\triangleright$ Chunk-and-Search
12:         ${\bm{Y}}^{r}=M_{l}^{\text{KI}}\left({\bm{X}},{\bm{H}},{\bm{M}}\right)$
13:         $\triangleright$ Knowledge Integration
14:         $\alpha=f\left({\bm{X}}\right)$
15:         ${\bm{Y}}=\alpha\cdot{\bm{Y}}^{m}+(1-\alpha){\bm{Y}}^{r}$
16:     else
17:         $({\bm{H}},{\bm{Y}})=M_{l}\left({\bm{X}},{\bm{H}}\right)$
18:     ${\bm{X}}={\bm{Y}}$
19:return ${\bm{Y}}$
@@ 3.1 Problem Formulation and Overview
Let $\mathcal{V}$ be a vocabulary, i.e. a set of discrete elements, of size $\left|\mathcal{V}\right|$. A model $M$ applies a function $f:\mathcal{V}^{*}\to\mathcal{V}^{*}$, taking as input a sequence of tokens from the vocabulary (of arbitrary length) while outputting a sequence of tokens from the vocabulary. We denote the input sequence as ${\bm{x}}=[x_{1},\dots,x_{T}$], which we refer to as the model’s prompt. The corresponding output sequence ${\bm{y}}=f({\bm{x}})$ is referred to as the model’s answer or generated response. Furthermore, a sequence-to-token mapping is a function $g:\mathcal{V}^{*}\to\mathcal{V}$ used to define $f$ through auto-regressive inference. Specifically, given an input sequence ${\bm{x}}\in\mathcal{V}^{*}$, the output tokens are generated one at a time using the recurrence: $x_{i+j}=g(x_{1},\dots,x_{i+j-1})$ and $f\left({\bm{x}}_{1:i}\right)=(x_{i+1},x_{i+2},\dots)$, where $1\leq i\leq T$ and $j\in\mathbb{N}$.
An LRM is defined by a state update rule $u:\mathcal{S}\times\mathcal{V}\to\mathcal{S}$ and an output function $r:\mathcal{S}\to\mathcal{V}$, where $\mathcal{S}$ is a finite set of states and a state is a representation of the system after processing a sequence from $\mathcal{V}^{*}$. Let $s_{0}\in\mathcal{S}$ be some initial state. Given some sequence ${\bm{x}}$ of length $L$, the state of the model at iteration $i$ is denoted by $S_{i}(x_{1},\dots,x_{i})$ and the output token is denoted by $R_{i}(x_{1},\dots,x_{i})$. These are defined recursively as:
@@
    $y_{i}=R_{i}(x_{1},\dots,x_{i})=r\left(S_{i}(x_{1},\dots,x_{i})\right)$.
Information from ${\bm{x}}$ flows to ${\bm{y}}$ through a state ${\bm{h}}\in\mathbb{R}^{d_{h}}$ where $d_{h}$ is fixed and finite. Thus for increasing sequence length or information dense settings, LRMs can struggle from the limited size of ${\bm{h}}$.
We observe that the LRM is directly limited by the size of its hidden state, which can be insufficient for modeling problems with many possible states, unless the hidden size grows with the size of the possible set of states. Resona introduces two flexible components to overcome this constraint without sacrificing the primary benefits of LRMs (namely parallel training and inference time efficiency): 1) a contextual search operation that operates on the input to retrieve context-specific information and 2) a knowledge integration component that mixes the retrieved information with the LRM output.
@@
2:Input: Input ${\bm{X}}\in\mathbb{R}^{T\times D}$ and LRM hidden state ${\bm{H}}\in\mathbb{R}^{T\times H}$.
3:Output: Attention mask ${\bm{M}}\in\mathbb{R}^{T\times T}$
[…]
6:Use $\mathcal{Q}$ to encode ${\bm{H}}$ into query embeddings $\bar{{\bm{Q}}}\in\mathbb{R}^{T\times E}$.
7:Compute chunk index sets for each $\bar{{\bm{Q}}}_{j}$: $\{I_{j}\}_{j=1}^{T}=\{\text{Top-}k(\bar{{\bm{Q}}}_{j},\bar{{\bm{C}}})\}_{j=1}^{T}$
8:With $\{I_{j}\}_{j=1}^{T}$, compute a mask ${\bm{M}}\in\mathbb{R}^{T\times T}$ such that ${\bm{M}}_{ji}=1\iff\left(i\in\{I_{j}\}\right)$.
@@
[…]
@@
$\alpha$ can be computed on an input-dependent basis for each element of ${\bm{Y}}$ or can alternatively be set as fixed hyper-parameter value. Because only the chunk-and-search design, each query attends to at most $kU$ elements from the initial input in $k$ contiguous blocks, enabling us to compute ${\bm{Y}}^{r}$ efficiently using existing sparse attention mechanisms. For simplicity, we maintain the use of a fixed constant $\alpha$ for the experiments that follow.
@@
2:Input: Input ${\bm{X}}\in\mathbb{R}^{T\times D}$ and LRM hidden state ${\bm{h}}\in\mathbb{R}^{T\times H}$, mask ${\bm{M}}\in\mathbb{R}^{T\times T}$
3:Output: ${\bm{Y}}^{r}\in\mathbb{R}^{T\times D}$
4:From ${\bm{h}}$, compute queries ${\bm{Q}}\in\mathbb{R}^{T\times d_{k}}$ using $W_{Q}$. In parallel, compute ${\bm{K}},{\bm{V}}\in\mathbb{R}^{T\times d_{k}}$ from ${\bm{X}}$ with $W_{K}$, $W_{V}$.
5:Compute multi-head attention output ${\bm{O}}=\mathrm{CrossAttn}\left({\bm{Q}},{\bm{K}},{\bm{V}},{\bm{M}}\right)$, where ${\bm{O}}\in\mathbb{R}^{T\times d_{k}}$.
6:Project ${\bm{O}}$ using $W_{\text{out}}$ into ${\bm{Y}}^{r}\in\mathbb{R}^{T\times D}$.